40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 3 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  2509   Mon Jan 13 19:04:38 2020 anchalDailyProgressFSSFSS Diagnostics - SN009 (North RFPD) Notch Improvement Attempt

Circuit changes (Updated Schematic)

I made the following changes to SN009:

  • R15 -> 10k Ohms for increasing Test Input Power.
  • R1   ->  226 Ohms
  • R2   ->  25 Ohms (Two 50 Ohm resistors stacked in parallel)
  • Added Ln, 2.2 uH fixed inductor.
  • Added Cn 0.5-1.5 pF Trim Tunable Capactior.
  • Ln and Cn are in series such that Ln + Cn is parallel to R1.
  • Rewired the 5V line from one end of C3 to one end of C14 (U8-14) with yellow 20 AWG single core wire.

Updated Photo


Measurements

  • I used the setup in 40m to take light transfer function of the modified SN009.
  • Method described in CTN:2247 was used.

Inference

  • It turns out, however, that the placement of "active" notch didn't help at all. Infact, the ratio between peak and notch went from 30 to 20 :(
  • I needed to reduce the values of R1 and R2 to ensure the notch doesn't effect the peak transimpedance much.
  • During the tuning, I found that as I turned the screw on the capacitor, the transimpedance will just cycle through phases of low and high impedance. I dont' think I ever saw a very clear notch because of the new changes.
  • I also tweaked the passive notch slightly to bring it closer to 72 MHz.
  • I'll think more about what has happened and will try another pair of inductor and capacitor.

Data

Attachment 2: SN009_D980454-00-C.pdf
SN009_D980454-00-C.pdf SN009_D980454-00-C.pdf
Attachment 3: SN009_Notch_Improvement_Attempt.pdf
SN009_Notch_Improvement_Attempt.pdf
  2508   Thu Jan 9 19:00:19 2020 anchalDailyProgressFSSTTFSS OLTFs

I repeated this measurement to compare later after changes to RFPDs.


Data

Attachment 1: FSS_OLTF.pdf
FSS_OLTF.pdf FSS_OLTF.pdf
  2507   Tue Jan 7 17:59:20 2020 anchalDailyProgressFSSNPMC TF

Method:

  • I injected excitation at FP1Test and readout at Mixer Out (FP3Test).
  • All TF were taken in three stages encompassing the whole frequency range. The same configuration files were used.
  • First, I took TF measurement when the loop was closed and the variable gain is 6 dB.
  • Then I opened the loop by removing mixed down PD signal from FP2Test and shorting it. This gave direct transfer function from FP1Test to Mixer Out as in second plot.
  • Using the two measurements, I calculated the open-loop transfer function for the NPMC Servo Loop.

Inference:

  • My hope was to see some activity or unity gain around 433 kHz where we see the spike of noise.
  • But the unity gain is much before around 6 kHz with phase margin of about 90 degrees.
  • There is a small dip around 433 kHz but again it seems like a reaction to the problem rather than the source of the problem.
  • The measurement, however, is pretty trashy. I could have used SR785 for the lower frequency region to make it better.
  • I think we are back to square one. Any comments to infer more or suggest further tests are welcome.

Data and Analysis

Attachment 1: NPMC_TF.pdf
NPMC_TF.pdf NPMC_TF.pdf NPMC_TF.pdf
  2506   Mon Dec 30 10:33:33 2019 anchalSummaryOtherSummary of questions asked in December

For the convenience of others, I'm summarizing the open questions I asked on elog in December. Comments on the posts, advice or answers to my questions would be nice.


  • CTN:2492 Why are output powers of these lasers much lower than rated power when diode current is maximum? And why do I see a maxima in South Laser Power when diode current is increased, shouldn't it just increase with diode current monotonically?
  • CTN:2495 Comments on beatnote noise differences with different laser power and modulation depth. Can I infer some good choice of laser power and modulation index from this? Suggestions for more tests on this line?
  • CTN:2496 Rana asked for the latest available photos of FSS RFPDS. We want to add an active notch in the final RF amplifier stage.
  • CTN:2497 There is indeed actual DC offsets at the output of FSS RFPDs RF out. Please cross-verify my measurement method and does this mean the MAX4107s are busted. Should I replace them?
  • CTN:2499 Time series measurement of the signal before and after the summing stage in FSS on Northside. There is significant leftover 1-Omega frequency even after the elliptical filter. I think I need to make the input end lossy to absorb back reflections from the elliptical filter. Need permission to modify the circuit and test if this helps.
  • CTN:2501 There is a logged event of South Laser losing about 20% of its output power with the same diode current. Is this laser dying? What is happening?
  • CTN:2505 As instructed, I have taken transfer functions through FSS from PZT path and EOm path to RIN before and after the PMC. I need validation on method used. My conclusion is the noise in FSS is due to RIN and not the other way round (FSS causing RIN).  I tried taking an OLTF of the PMC loop but the NPMC Loop unlocks as soon as I connect the source port of AG4395A to an excitation port in a summing stage. So I've been unable to verify the source of this RIN yet, but previous measurement (CTN:2502) suggests it is PMC.

HAPPY HOLIDAYS

  2505   Fri Dec 20 18:08:44 2019 anchalDailyProgressFSSPZT, EOM to RIN TF Measurement Through FSS

TF Measurement Method

  • I switched on Test1 switch on RF Board to cut off RFPD input to FSS.
  • I switched on Test2 switch to connect the excitation port at the summing point.
  • I switched on Ramp Engage for FSS to ensure paths to PZT and EOM are connected.
  • Then, I disconnected and shorted EOM input when taking the measurement for PZT and vice versa.
  • COM Gain was 6dB and FAST gain was 10 dB.
  • TF was measured from Test2 EXC Port to the output of PDs. This was divided by 10s averaged DC value of PD output to get TF to RIN.
  • Measurements were taken in turns for low and high frequencies with IFBW 10Hz and 1 kHz.
  • Again, PDs are  Thorlabs PDA10CS placed at (110,44) at dumped end of a PBS before PMC and at (64,28) at dumped end of Input PBS of Faraday Isolator after PMC.

Calculations:

  • The noise spectrum in FSS loop was measured at OUT2 port on the common board.
  • This was divided by 50/(453+50) voltage division caused by 50 Ohm input impedance of AG4395A. This gives spectrum at the output of the U3 adder stage in the common board.
  • The TF measured above are divided by 392/1200 which is the gain from Test2 EXC to output of U3. So the TFs are no referenced from the output of the U3.
  • The measured noise spectrum is multiplied by the TF from that point to RIN at After and Before PMC due to PZT or EOM.
  • These are summed quadratically to get total estimated RIN ASD (before and after PMC) due to the noise from FSS.
  • Finally, I also plotted the RIN ASD measured in CTN:2502 before and after PMC when FSS was ON.

Inferences:

  • The measured RIN, both after and before PMC are much higher than the estimated noise due to FSS.
  • In particular, the bump around 10kHz in measured RIN before and after PMC is not present in transfer functions from FSS to RIN or the noise spectrum itself in FSS.
  • The notorious 433 kHz peak is present everywhere, but there's order's of the magnitude of difference between FSS contribution and seen RIN. So it is mostly the other way round, the RIN causing these peaks in FSS.
  • Either that, or there is something wrong with the way I took measurements or did the above calculations.
  • Also, I'm aware that the RIN is actually fed back into the FSS through PD and it is not an open-loop. But I did this simpler analysis first. Maybe that's what is wrong.

Data

Edited on Mon Dec 23 15:21:50 2019 .

Attachment 1: NorthLaserPZTEOMtoRINTFThruFSS.pdf
NorthLaserPZTEOMtoRINTFThruFSS.pdf NorthLaserPZTEOMtoRINTFThruFSS.pdf NorthLaserPZTEOMtoRINTFThruFSS.pdf
  2504   Fri Dec 20 15:00:30 2019 anchalDailyProgressFSSPZT, EOM to RIN TF Measurement

I took some TF measurements, but I'm not sure if I used the right method to do this. All the resutls look essentially the same, so maybe I'm just measuring instrument noise and nothing else. Regardless, I' posting the results. 

Measurement details:

For both measurements, I sent the source signal, first directly to PZT or EOM and second with a 50 Ohm termination in parallel. In case of EOM, I also did another measurement with low power to see if I can uncover any saturation effects my high powered source might be causing.

The output of two photodiodes, Thorlabs PDA10CS at (110,44) at dumped end of a PBS before PMC and at (64,28) at dumbed end of INput PBS of Faraday Isolator after PMC was measured and I took DC level of the outputs averaged over 10 s right after taking TF measurement. This was divided by the measured TF to get units of 1/V i.e. from PZT/EOM to RIN.

The parameter configuration files for the measurements are in the data directory.

What might have gone wrong:

I suspect that maybe AG4395A is unable to drive the capacitive load of PZT and EOM after a certain frequency. I need to find a better way to actuate the PZT and EOM in a known fashion, possibly through the FSS box or some buffer driver.

In the case of EOM, I'm also uncertain if the amplitude of actuation would be enough to do anything whatsover. Maybe the transfer function needs to be taken with high voltage driver.

Any comments on my measurement techniques are welcome as I'm surely not doing this right.

Rana's question about PMC:

I concluded that PMC is causing the 433 kHz peak because the only time I do not see it in CTN/:2502 is behind PMC when FSS is OFF. I couldn't think of a way to check if PMC Servo is causing it on its own. Can I do that without closing the loop somehow?


Data

Attachment 1: NorthLaserPZTEOMtoRINTF.pdf
NorthLaserPZTEOMtoRINTF.pdf NorthLaserPZTEOMtoRINTF.pdf NorthLaserPZTEOMtoRINTF.pdf
  2503   Wed Dec 18 22:03:17 2019 ranaDailyProgressFSSMaybe the NPRO Laser isn't bad after all

Could be that either

  1. the FSS makes RIN through driving the PZT / making beam jitter
  2. FSS drives EOM and makes polarization rotation if EOM is misaligned or beam is not linearly pol and aligned to EOM axis

I suggest measuring the RIN/PZT and RIN/EOM TFs, both before and after the PMC.

Is the PMC servo making the 430 kHz peak?

  2502   Wed Dec 18 18:40:32 2019 anchalDailyProgressFSSMaybe the NPRO Laser isn't bad after all

Today I inserted two photodiodes, Thorlabs PDA10CS at (110,44) at dumped end of a PBS before PMC and at (64,28) at dumbed end of INput PBS of Faraday Isolator after PMC. These photodiodes were set at 20dB gain which according to the manufacturer gives us a bandwidth of 1.9 MHz.

Then, I just took Noise Spectrums using AG4395A with the parameter file attached in the data directory, from 0 to 1 MHz at 1 kHz ResBW and 50 averages. The output was divided by the DC level of the photodiode which was measured using TDS 3052B averaged over 10s. This measurement was done with FSS ON or OFF.


Inference

  • Whenever FSS is ON, the loop is completed, so noise due to PMC and FSS is present in the measured RIN.
  • But when FSS is OFF, before PMC, the noise should be just of free running laser even though PMC is locked downstream.
  • The most important observation is that when FSS is OFF, before PMC, the RIN is pretty small with no peaks in 100s kHz, particularly the notorious 433 kHz peak.
  • The dome from 1kHz to almost 100kHz peaks at 10kHz is most probably due to FSS as it goes away whenever FSS is OFF.
  • The peak at 433 kHz persists even after switching off FSS and is present only after PMC. So this peak is clearly due to PMC and not due to laser as was thought before.
  • Conclusion is, that laser is free of any large RIN and most of the RIN introduces is due to FSS or PMC. 
  • I've left the PDs in there with termination. I'll change some PMC loop and FSS loop parameters tomorrow to see nature of their contribution to RIN.

North Laser is hereby acquitted.


Data

 

Attachment 1: NorthLaserRINAnalysis.pdf
NorthLaserRINAnalysis.pdf NorthLaserRINAnalysis.pdf NorthLaserRINAnalysis.pdf
  2501   Wed Dec 18 14:04:43 2019 anchalDailyProgressLaserSouth Laser Power Dropped.

After the last calibration, South Laser power at Laser head determined by its internal photodiode and show as setting PWR today (noticed today) dropped to 80 mW with none of the other settings changed. The diode current is still the same, so this could be one of the following two things:
1) The internal circuit to measure laser power at head went wrong.

2) Something is wrong with the laser crystal.

The laser power as seen by the reflection photodiode at South PMC captured a glitch on Dec 17th, 2019 8:30 am (might be 7:30 am in PST). The glitch shows that PMC went out of lock due to it but then returned to a laser power lower than before. The voltage level decreased from 2.72 V to 2.3 V after lock, which is equivalent to a drop of ~15.4%. The laser head power level according to PWR monitor dropped from 101 mW to 80 mW, which is equivalent to ~20.8%. While these don't match, PMC reflection is also not a true measure of power level. But I am sure this power has actually reduced and even if the laser head power meter is off a little bit, we have witnessed a big drop in power.

Restarting the laser (Put it on Standby and switch it back on) didn't change the power level.

Restarting the laser with turning the key off and on also didn't change the power level.

Even though 80 mW is enough for our experiment, this sudden decline in power shows there is something happening with the laser that we do not understand.


Data

Attachment 1: SouthLaserPowerDropped.pdf
SouthLaserPowerDropped.pdf
  2500   Tue Dec 17 18:29:58 2019 anchalDailyProgressFSSNeed for ISS?

If we are going to keep the same laser with busted Noise Eater, should be go in the direction of implementing the ISS?

In Oct, 2018, I along with Johannes developed ISS boards and photodiode transimpedance amplifier boards. These were characterized by Johannes at Cryo_Lab:2180 and Cryo_Lab:2181. However, Johannes was using AOM while we have EOAM here. There are slight differences which should be addressable given we have an offset port also in the ISS board. But the expected open loop gain needs to be worked out along with right choice of transimpedance and cavity pole neutralizer stage in ISS.

Johannes gave me his M2ISS photodiode mounts before leaving, so we have them too. This looks like a 1-week project from start to finish. So with the thumb rule of "nothing goes as expected" and me being a grad student, this should take 2 weeks. Is it worth it? Should I start working on this? Or can I just get a laser with working NE?

Also, there is no real documentation of why we stopped using the existing ISS in CTN? Last mention of ISS is at CTN:2132 and after that I came and was working on ISS boards which we never installed in CTN since or FSS, RFPDs and thermal controls were priority then.

  2499   Tue Dec 17 16:55:07 2019 anchalDailyProgressFSSFSS Diagnostics - TTFSS North U7 Boost Stage Input-Output TimeSeries

As suggested few group meetings ago, I took time series data at the input and output of boost stage opamp U7 (at TP15 and TP16) using TDS 3034C. I know trusting oscilloscope for synchronous measurement of two channels is not a great idea, but this is a good zeroth order approximation for this approach. 500 MHz 10MOhm impedance probes were used for the measurement of the signals. The data is taken at two different acquisition rates and I used basic math to calculate the expected ideal output. Following formula was used:

V_{Ideal}(t) = V_{TP16}(0) \ + \ V_{TP15}(0)\frac{R_{29}}{R_{37}+R_{39}} \ - \ V_{TP15}(t)\frac{R_{29}}{R_{37}+R_{39}} \ -\ \frac{1}{C_{35}(R_{37}+R_{39})}\int_0^tV_{TP15}(t)dt

Ofcourse the offset point above wouldn't work. Also, I have not included the notch filter that is also present at this stage.

The results of this analysis are attached. I also, calculated the required slew rate of the AD847 at this position. At the 100 MSa/s sampling rate, we saw that maximum required slew rate for ideal signal was around 110 V/us, much less than rated 300 V/us. I don't see a bad change in shape of the waveform. At 1 GSa/s sampling rate, we see that  required slew rate reached about 220 V/us at indeed oscillations at high frequencies are limited even though this is below the rated value.

Looking from data taken at CTN:2474 present in this directory, these fast oscillations that we are seeing on the oscilloscope are around the modulation frequency 36 MHz.  These leftover downconverted 2-Omega from teh RFPD might be saturating slew rate limtis of many opamps in the TTFSS box. I think I should look back into Andrew's implimentation of the elliptical filter right after the mixer in the RF board. He did say that we need to make the filter lossy to ensure the reflected signal at 36 MHz gets absorbed at a 50 Ohm resistor. For this we need to add a resistor to gnd at node rfbn2. Seeking permission to make modifications to board.


Data

Attachment 1: NFSS_UT_TS_Analysis.pdf
NFSS_UT_TS_Analysis.pdf NFSS_UT_TS_Analysis.pdf NFSS_UT_TS_Analysis.pdf NFSS_UT_TS_Analysis.pdf NFSS_UT_TS_Analysis.pdf NFSS_UT_TS_Analysis.pdf
  2498   Mon Dec 16 09:43:56 2019 Ian MacMillanNotesFSSFSS Plant Model v2

In talks with Anchal, we have adapted our model and the transfer function fits rather well. These measurements were taken with the AGmeasure and SRmeasure script on the larger instruments in the CTN lab (I can't remember the exact make of them). Two were used for different frequency bands: one up to 100 kHz and the other one above. In order to have continuous measurements, the high-frequency measurement was shifted to be continuous with the lower frequency measurement. This was not a significant shift and does not impact the overall interpretation. It is merely a correction for different calibrations in the Instruments used to measure the transfer functions.

The Magnitude of the transfer functions matches well and Anchal agrees that they also look good. So the circuit can be used in its current form. It isn't currently in a box but that is an easy fix. I am still making full documentation on it and that should be done over break. I am also looking to make it as a single PCB to eliminate any points of failure that come from attaching multiple PCBs together. It is not entirely necessary because as is seen in the graphs, the circuit works well but designing a PCB is a good task for over break because it does not really require any in lab work and will be highly beneficial down the road. I am still talking to Anchal about the circuit and its implementation.

Attachment 1: PZT_TF.pdf
PZT_TF.pdf
Attachment 2: EOM_TF.pdf
EOM_TF.pdf
Attachment 3: Non-Moku_Data_PlantCir_V1.zip
  2497   Fri Dec 13 17:44:05 2019 anchalDailyProgressFSSFSS Diagnostics - RFPD RF Coupled Output Offset

I rechecked the 20dB coupled RF output of SN009 (RFPD on north Cavity reflection) and SN010 (RFPD on south Cavity refleciton). The following are the mean values over 10s taken at DC coupling with 50 Ohm input impedance, +/- 40 mV input range with two different oscilloscopes.

SN009 RF Out 20dB Coupled
  Shorted SN009 RF Out 20dB Coupled SN010 RF Out 20dB Coupled
TDS 3052B -0.2 +/- 0.01 mV 7.0 +/- 0.1 mV 7.3 +/- 0.1 mV
TDS 3034C 0.5 +/- 0.01 mV 7.4 +/- 0.1 mV 7.6 +/- 0.1 mV

 

 

 

 

Note that these are 20dB coupled values, so the actual offset reaching the FSS board is about 70 mV. The PD input is AC coupled through a transformer so it shouldn't be reaching further than there but ideally MAX4107 at the RFPD is supposed to have a maximum input offset voltage of 3 mV which at gain 10  should look like 30 mV. So what we are seeing (7.0 mV) is more than twice the rated maximum input offset. I'm not sure if this means our MAX4017 is busted or something is wrong in the loop. Help needed in understanding this result.

  2496   Fri Dec 13 16:24:02 2019 anchalDailyProgressFSSFSS Diagnostics - RFPD Circuit Latest Available Photos

I found the latest available photos for the Cavity Reflection RFPD circuits.

SN009 as of March 25, 2019
SN009 as of March 25, 2019
SN010 as of April 1st, 2019
SN010 as of April 1st, 2019

There were minor changes made after taking these photos which are logged at: https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:experiments:psl:rfpd

I'll update these above photos whenever the next time I get a chance. They would be present at:

SN009: https://git.ligo.org/cit-ctnlab/ctn_electronics/blob/master/RFPD/Photos/SN009.jpg

SN010: https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:experiments:psl:rfpd

Attachment 3: SN009.jpg
SN009.jpg
Attachment 4: SN010.jpg
SN010.jpg
  2495   Thu Dec 12 19:15:55 2019 anchalDailyProgressFSSFSS Diagnostics - Beatnote Spectrum wrt Laser Power and Modulation Index

I took beatnote spectrums in the current modulation index (was set to around 0.3 earlier). Then I took spectrum after attenuating the modulation signal power to both EOMs on North and South Path by 3 dB and 6 dB. This should reduce the modulation depth by 1/\sqrt{2}} and 1/2. After every change, a small displacement happens in the thermal control of the cavities, so I had to wait for some time to let it settle. The gains of the FSS loops were kept constant to make sure only the modulation depth is the parameter that is changing. Gain values were 24 dB and 16 dB for South COM and FAST gains, and 11 dB and 10 dB for North COM and FAST gain.

Inference:

  • There is almost no difference in noise upto1 kHz.
  • There is a new bump I'm seeing between 100 and 200 Hz though which was there on Dec 12th but not on Dec 13th measurements.
  • This bump might have been reduced due to a decrease in laser power or something might have changed between the two days. I'll check with increasing power what happens.
  • Above 1 kHz, we see that as modulation depth is deceased, we see a reduction in noise. The same goes for power.
  • Overall, this does show that the effect of these changes is more or less as expected and the effect of non-linearity, if any is not much.
  • It is still valuable to quantify properly how much the non-linearity might be affecting the final beatnote spectrum though. Working on it.

Data

Attachment 1: BeatnoteSpectrumwrtPowerandModindex.pdf
BeatnoteSpectrumwrtPowerandModindex.pdf
  2494   Thu Dec 12 19:05:33 2019 ranaNotesFSSFSS Plant Model v2

please overlay all the plots and also include the ideal TFs - its hard to understand anything without the comparisons

  2493   Thu Dec 12 11:54:11 2019 Ian MacMillanNotesFSSFSS Plant Model v2

Testing the circuit using SRmeasure and AGmeasure, I got more accurate readings of the transfer functions. The PZT path TF looks good. The EOM path looks good up until 105 Hz, then it doesn't drop like it is supposed to. My next task is to fix the high frequency EOM path. These graphs are a work in progress. I will include plots with the ideal TF at a later time.

Attachment 1: PZT_LowFq_V1.pdf
PZT_LowFq_V1.pdf
Attachment 2: PZT_HighFq_V1.pdf
PZT_HighFq_V1.pdf
Attachment 3: EOM_LowFq_V1.pdf
EOM_LowFq_V1.pdf
Attachment 4: EOM_HighFq_V1.pdf
EOM_HighFq_V1.pdf
  2492   Tue Dec 10 17:03:41 2019 anchalDailyProgressLaserLaser Settings back to defaults

I put laser settings on both North and South Cavities back to default. From this point onwards, all settings about the lasers would be known and kept track of. The red values are the settings that were changed.

NPRO Laser Settings

Property Display Symbol North South Units Notes
Laser Model - M126N-1064-700, SN 5519, Dec 2006 126N-1064-500, SN 280, Nov 1997 -  
Diode Temperature DT 22.3 28.7 ^\circ C Informational only.
Diode TEC Voltage DTEC 0.8 0.7 V Informational only. +ve -> cooling, -ve -> heating.
Measured Laser Crystal Temperature LT 40.8 55.2 ^\circ C Informational only. Calibration dependent.
Laser TEC Voltage LTEC 0.0 -0.5 V Informational only. +ve -> cooling, -ve -> heating. Manual says typically should be 0.0V.
Target Laser Crystal Temperature T 40.087 -> 40.0010 48.0010 ^\circ C Changed back to factory set value on North Side.
Laser Head Power Level PWR 66  ->  624 92  ->  101 mW Calibration dependent. CHanged the calibration to meet the power meter but even then, power meter says a maximum 500 mW, so North Side is not entirely correct. On the south side, it was difficult to mount power meter perpendicular to the beam, so there might be some clipping loss in calibration.
Power Adjustment ADJ 0 -2  ->  0 - From -50 (off) to +10. Changed the diode current around set value.
Diode Current DC 2.06 2.04 A It can be changed to change power level. Reflects measured value.
Diode Power Monitor DPM 0.00 0.00 V

Calibration Dependent.

Noise Easter NE ON ON - -
Laser Diode Status LD ON ON -  
Nominal Diode Current - All the way clockwise All the way clockwise - It can be changed by turning the left potentiometer from the back of the laser head. Factory default is all the way clockwise. I have set both North and South Lasers to this point.

Note:

While turning the nominal diode current of south laser all the way clockwise, I found that the laser power peaks before the maximum diode current is reached. This diode current is about 1.9 A. This is unexpected. Any explanations on this would be helpful.

  2491   Tue Dec 10 11:57:58 2019 Ian MacMillanNotesFSSFSS Plant Model v2

 

First, off I forgot a resistor on the previous diagram so here is the fixed version. Also, Anchal said that the EOM noise is probably caused by instrument noise at the low end. The phase response should be flat on both and on the PZT path the non-flat response is caused by some high pass filter caused by AC coupling or something of the sort.

I am still trying to figure out how to get a flat response on the phase. Anchal did some math that looks at the possibility of it and I am going through it.

Attachment 1: IMG_8124.jpeg
IMG_8124.jpeg
  2490   Mon Dec 9 19:54:44 2019 anchalDailyProgressFSSFSS Diagnostics - How much distortion affects the functioning of PDH?

I'm trying to think hard with my small brain how the distortion would affect the PDH functioning and inject noise in the frequency of the laser. I have a line of reasoning which starts with a question.:

  • The two RF sidebands of laser that fall on the cavity, upon reflection, do they destructively interfere? Because the calculation I did in CTN:2481 suggests that the power at second harmonic (2-Omega frequency) is so high because of the sidebands beating with each other.
  • In case they do not destructively interfere, the generated second-order harmonic (2-Omega) will have laser amplitude noise on it. This when it mixes with the (1-Omega) signal which actually carries the cavity length noise at the inputs of MAX4107, the distorted 1-Omega created would have laser amplitude noise on it.
  • This might be the reason why FSS loops get overwhelmed with seeing a lot of laser amplitude noise which ideally it is not supposed to see. And it tries to correct this noise in phase quadrature making the situation worse.
  • Since Andrew left, I increased the laser power reaching the cavities to go above the shot-noise limit of the photodiodes. Maybe, this increased light level increased this effect to a point where we are witnessing the problem.

Of course, all this depends on the RF sidebands interfering constructively upon reflection. I remember (I don't know from where) that it is the opposite. Either there is a fault in my calculations or this is indeed what is happening. I need to understand this properly to go further. Need help.

  2489   Mon Dec 9 18:08:13 2019 Ian MacMillanNotesFSSFSS Plant Model v2

I tested the circuit and got the attached results. The PZT transfer function looks reasonable. some adjustments to the values of the resistors in the circuit should have the match up. The phase is still out of alignment for both. The transfer function of the EOM looks far from the model. I suspect that this is due to a bad solder point or poor choice of components rather than the circuit's design.

Attachment 1: OutputGraphsTransferFunction.pdf
OutputGraphsTransferFunction.pdf OutputGraphsTransferFunction.pdf
Attachment 2: _Circuit_Design_V1.zip
  2488   Thu Dec 5 18:25:11 2019 ranaDailyProgressFSSFree running laser frequency noise spectrum

when the NPRO crystal is oscillatin

          no tellin what is relaxin

  2487   Wed Dec 4 18:11:09 2019 anchalDailyProgressFSSFree running laser frequency noise spectrum

I took a spectrum of PMC error signal when the FSS loop is not closed. This should provide a rough estimate of the free running laser noise. We had earlier seen a peak at 435 kHz in the Northside, hence I wanted to take this data with some references. First of all, this peak is very similar in the description of relaxation-oscillation peaks of these NPRO lasers mentioned on page 52 of this manual. The "Noise Eater (NE)" is supposed to suppress this peak significantly. However, in the spectrum of the PMC error signal, there is no difference when noise eater was ON or OFF.

I took a spectrum of Southside as well, just to see if I could see action of Noise eater there. For south laser, the noise eater suppressed noise only till 100 kHz or so and probably this side also has a similar relaxation-oscillation peak problem but is shadowed by a large feature at 30 kHz. Not, the absolute value of the spectrum between North and south are vastly different due to different amount of light, different transimpedances od the PDs and different gain values in the feedback circuit.


However, the noise eater is supposed to reduce relative intensity noise only. And the error signals of PMC should really be telling us noise in the frequency of the laser. So maybe I'm connecting two dots in different Hilbert spaces. But Rana suggested that a busted Noise Eater could be the reason for the 435 kHz peak, I just do not understand how RIN would cause frequency noise so badly. I thought photothermal transfer functions from RIN to frequency noise were very small.

Attachment 1: PMCErrorSignals.pdf
PMCErrorSignals.pdf PMCErrorSignals.pdf
  2486   Wed Dec 4 17:09:44 2019 anchalDailyProgressTempCtrlIncreased range of out-of-loop temperature sensor

This has happened few times now that acromag channel for the can heater driver stopped updating according to the PID script and the can gets heated to a very high temperature. This pushes the temperature out of the ranges of the current AD590 temperature sensor board. I have changed the range of channel 2 (this was being used for out-of-loop) to ensure we can still see some meaningful temperature value when such incidents happen. I have replaced R18 from 100k to 27k. The updated table is:

CH No on Board EPICS Channel Name Temperature Conversion Function (ºC) Range (ºC)
1 C3:PSL-TEMP_TABLE = V/0.810875 + 27.30248869 13.860-40.745
2 C3:PSL-TEMP_VACCAN_OOL = V/0.43875 + 33.3115385 8.450-58.155
3 C3:PSL-TEMP_VACCAN_INLOOP = V/1.625 + 33.3115385 26.604-40.020

Weird phenomenon?

  • I'm not sure this problem occurs though. Right now the out-of-loop temperature sensor shows that the temperature of the can is 55 Degree Celsius.
  • This is also not cooling down fast. Last time it took days for it to cool down to set value.
  • The alignment of the cavities change vertically when the can temperature is significantly different from the set value of 34.38 Degree Celcius.
  • This hinders them from locking properly unless I tune the alignment back. But when the can will cool down finally, they will be misaligned again.
  • Also, I just refuse to believe it is actually that hot, but I check the voltage after the transimpedance amplifier in temperature sensor boards and two independent AD590s are reporting this.
  • And anyways, why is the acromag output channel getting frozen anyways.
  • That being said, this is an irreproducible but non-harmful problem yet, so lower priority than the FSS saga going on.
  2485   Tue Dec 3 17:27:54 2019 Ian MacMillanNotesFSSFSS Plant Model v2

The new circuit design has been built and is shown below. I tested the transfer functions and the PZT path looks good by eye but the EOM path looks like there is a solder point that is bad. I will show the graphs when I have them done.

 

Quote:

With discussions with Anchal and some reading

It seems impossible to create a casual circuit with a zero phase shift. (See this for more) 

If we have a circuit with an impulse response h(t) and transfer function H(f)=F[h(t)] where H(-f)=H*(f). For the filter to cause no phase shift then ∠H(f)=0 for a complex exponential input for all f. It is also impossible to have a constant phase shift unless that phase shift is zero.

Therefore "filter does not change the phase at all, then H(f) is a real-valued function, and because of the conjugacy constraint, it is also an even function of f. But then its Fourier transform h(t) is an even function of time, and thus the filter cannot be causal (except in trivial cases): if its impulse response is nonzero for any particular t>0, then it is also nonzero for −t (where −t<0)" 

Because this can't be done casually, it should be done using a Field Programmable Gate Array. Unfortunately, I don't think we have access to one. I am reading up on the Moku FIR Filter builder to find out if we can program it to do what we want.

Quote:

I have designed a passive circuit that seems to match the ideal transfer functions in shape. Scaling should just be a game of playing with the values of the resistors and capacitors. The phase still seems to be an issue. There is an unwanted phase shift from 0 -> -90.

The next step is trying to finalize the values for the resistors and caps. Possibly model in zero if I have time. Then build and test. Also fix the phase.

Quote:

I have updated the plant model to contain the cavity pole also. Cavity pole is a pair of positive and negative real poles, so it is hard (or maybe impossible) to imitate it exactly with an electronic circuit. Or maybe, my analysis is wrong.

Nevertheless, I have for now made this circuit which has a second-order pole, so it correctly matches the magnitude of the model transfer function up to 1 MHz for both PZT and EOM paths. Note that the elliptical filter is not included in this as we can connect the circuit to Test port 1 which injects just before the filter in LIGO-D0901894. Also, for the gains in EOM path, I had to add some factors to make it the same as the model transfer function. All components are calculated for E12 series resistors and capacitors.

Attached is a pdf of the notebook which contains all the mathematics in latex and a zip file with all files to recreate and further work on this. Ian can use these as support to learn zero further.

 

 

 

Attachment 1: IMG_8065.jpeg
IMG_8065.jpeg
Attachment 2: IMG_8069.jpeg
IMG_8069.jpeg
  2484   Wed Nov 27 16:31:20 2019 ranaDailyProgressFSSFSS Diagnostics - Quick HF OLTF of NFSS

I doubt it.

When feedback loops be noisy and busted,

the swept sine TFs cannot be trusted

Quote:

So, we can say with some confidence that the 435 kHz signal seen in the oscilloscope in CTN:2482 at TP1 is actually due to some non-linear effect most probably

 

  2483   Wed Nov 27 15:17:51 2019 anchalDailyProgressFSSFSS Diagnostics - Quick HF OLTF of NFSS

I quickly took a high-frequency Open Loop Gain measurement of NFSS loop at 10 dB COM Gain and 10 dB FAST gain, using the same measurement method as in CTN:2443. The UGF has not changed much but there is a dip at 435 kHz. This was there before too, I was just not paying enough attention to this part of OLTF before. So, we can say with some confidence that the 435 kHz signal seen in the oscilloscope in CTN:2482 at TP1 is actually due to some non-linear effect most probably and does not get suppressed at all. The phase margin near UGF looks about 135 degrees so there is no solid reason to believe this could be due to loop oscillation.

So I got to think of what combination of RF frequencies might be mixing down to create this oscillation and where. This oscillation is also visible in Plot 6 and 7 of NFSS_RFPD_Output_Oscilloscope.pdf of the measurements done in CTN:2470.


Data

Attachment 1: NFSS_OLTF_HF.pdf
NFSS_OLTF_HF.pdf
  2482   Wed Nov 27 13:59:24 2019 anchalDailyProgressFSSFSS Diagnostics - TTFSS TP1 TimeSeries

I took time-series data at TP1 on NFSS. This is just after the elliptical filter which is after the demodulation on board.

As also seen in the spectrum measurement at this testpoint in CTN:2474, there is a lot of power at around 435 kHz.  But this is not noise!

As seen on the oscilloscope, this is a near-perfect sinusoid. So this must be either of the following:

  • A mixed down frequency due to non-linearity of MAX4107 or the Mixer JMS1-H .
  • This could be an oscillation in the closed-loop of FSS at the unity gain frequency. But last I checked (CTN:2443), the UGF for the same COM Gain (10dB) and FAST Gain (10 dB) was 369 kHz.

This measurement was taken with a 500 MHz 10x Probe with a 300MHz TDS 3034C oscilloscope at 0.5 GSa/s sampling rate.

Interestingly, there is no such oscillation or peak on the South side. However, the south sides COM Gain is 24 dB and Fast Gain is 14 dB. So it could be because it is just suppressing this non-linear effect properly or just has a very high UGF.


Data

Attachment 1: NFSS_TP1_Oscilloscope.pdf
NFSS_TP1_Oscilloscope.pdf
  2481   Tue Nov 19 18:51:02 2019 anchalDailyProgressFSSFSS Diagnostics - Second-order Distortion Calculation

I did some theoretical calculations using the datasheet value of second harmonic SFDR from MAX4107 and the transfer function I measured from Test IN ports of RFPDs (using 100 kOhm series resistance).

Calculation:

  • Attached notebook has all the calculations. In all places, "Omega" refers to the modulation frequency.
  • I first calculated power at 0, 1-Omega and 2-Omega from the resonance frequency in the light incident on the cavity using Bessel functions.
  • I used reflection function from the cavity (using transmittance of 5 ppm, cavity length of 1.45" and assuming lossless cavity):
    \LARGE R = \frac{- r(1 - e^{\iota \omega/\nu})}{1 - r^2 e^{\iota \omega /\nu}}
  • Using this, I calculated reflected power at  0, 1-Omega and 2-Omega from the resonance frequency.
  • Then I used -40 dBc SFDR for second harmonic generation mentioned in MAX4107 datasheet when used with a gain of 10. I used following formula to calculate this:
    \LARGE P_{1\Omega Dist} = \frac{(V_{1\Omega}V_{2\Omega})}{100}10^{-\frac{SFDR}{10}}
  • Finally, I also calculated ratio of 1-Omega signal (actual beat between Sideband and Carrier) and the 1-Omega Distortion signal (generated) due to the non-linearity of MAX4107.

Inferences:

  • These are purely theoretical calculations but I can put in the second harmonic SFDR by measurement tomorrow.
  • The suppression of 2nd harmonics by the notch in the RFPD circuits is very bad. It only suppresses the second harmonic by 25-30 dB.
  • This is not enough as this 2-Omega signal would mix a lot with 1-Omega and would also go forward into the TTFSS box after downconversion.
  • The ratio of Signal at 1-Omega to the distortion at 1-Omega is about just 8 at 100 Hz. This doesn't reach 100 even till 1 kHz.
  • And below 10 Hz, the distortion dominates the 1-Omega signal from RF out.
  • This looks like a big limitation to our FSS loop. The notch filters in the RFPDs need to be doing much better job then they are doing.

Edited Wed Nov 20 14:43:07 2019: Corrected an error in code.

Attachment 1: FSSDistortionCalc.pdf
FSSDistortionCalc.pdf FSSDistortionCalc.pdf FSSDistortionCalc.pdf FSSDistortionCalc.pdf FSSDistortionCalc.pdf FSSDistortionCalc.pdf
Attachment 2: FSSDistortion.zip
  2480   Thu Nov 14 12:19:32 2019 ranaDailyProgressFSSFSS Diagnostics - Two Tone Third Order Intermodulation Test

cool - how bout use this new no-how to estimate the excess in FSS?

  2479   Tue Nov 12 18:36:59 2019 anchalDailyProgressFSSFSS Diagnostics - Two Tone Third Order Intermodulation Test

In line with industrial practices, I did two tone third order intermodulation test today on the FSS RFPDs. This test was inspired by procedure described in this technical note by MiniCircuits and this paper at IEEE.

Method

  • Similar to previous measurement, two RF tones are generated using Moku.
  • However, now, they are separated by an audio frequency, 10 kHz in our case. This is marked as DF in the plots.
  • These tones are centered around the resonant frequency of the RFPD marked as CF in the plots.
  • The two signals are combined together through ZFSC-2-1W+.pdf, but this time all ports of the combiner have an attenuator on it.
  • 3 dB attenuators at the input port and 6 dB at the output port. This was mentione din MiniCircuits as a step to improve impedance matching near the combiner.
  • Ideally, two low pass filters also should be put at the input port, but I couldn't find any with high enough cut-off frequency.
  • The combined source signal was getting attenuated by 8.75 dB. So I just drove the input signals higher by this amount.
  • The Source signal is fed to the Test IN of the RFPD and RF out is read through AG4395A.
  • The power in each harmonic is calculated from the spectrum and plotted together against source power.
  • Since all RF signals are the near-resonant peak of the photodiode, there would be little to no difference in TF seen by them.

Inference:

  • I think because of the 100 kOhm resistor at the Test In port of RFPDs makes the source signal too feeble.
  • I'm just measuring instrument noise upto source power of -10 dB. I can't go above 0 dB in current setup atleast.
  • At 0 dB source power, which is equivalent to 3 uApk photocurrent from the photodiode, the third-order harmonics are 60 dB lower.
  • One way of reporting this is, Spurious Free Dynamic Range (SFDR) at 36(37) MHz is 60 dBc at 0 dBm source power of SN009 (SN010).
  • Another way this figure is reported in industry is by extrapolating at which point the third-order harmonic power becomes the same as fundamental.
  • But I do not have enough distortion showing data points to extrapolate to calculate the intercept.
  • Last plot is when I connected the cables together which go to DUT. We see SFDR of 60 dBc here itself.
  • The last two data points are when input range of AG4395A are very high and hence has increased noise floor.
  • So probably in my measurements also, I'm just seeing distortion due to measurement apparatus itself.
  • But this at least sets an upper bound of distortion to SFDR of 60 dBc at the resonant frequencies (within the errors that I, a graduate student, can make).

Datasheet for MAX4107

  • The datasheeet says that between 30-40 MHz, the SFDR is approximately 50 dBc (this is actually lower than what we measured).
  • Also, the third-order intercept is mentioned at 15 dBm source power.

Data

Attachment 1: FSStwoToneThirdOrderIM.pdf
FSStwoToneThirdOrderIM.pdf FSStwoToneThirdOrderIM.pdf FSStwoToneThirdOrderIM.pdf
  2478   Tue Nov 12 11:39:48 2019 anchalDailyProgressFSSFSS Diagnostics - RFPD RF Ouput under inspection

I calculated these values by integrating in the 8 MHz neighborhood around the marked harmonic peak, the power spectral density using the frequency at the point as the lower edge of the bin. Slew rate is calculated by multiplying the rms voltage level with the frequency and the fraction is calculated against the datasheet value for Max 4107.

NFSS RFPD Output Slew Rate Usage (MAX 4107, SR: 500 V/us)
Freq (MHz) Vrms (mV) Required Slew Rate (V/us) Fraction of Slew Rate Used (%)

36

59.221
2.1319
0.4264
72
9.248
0.6659
0.1332
108
1.135
0.1226
0.0245
144
0.39
0.0562
0.0112
180
0.621
0.1117
0.0223

 

 

SFSS RFPD Output Slew Rate Usage (MAX 4107, SR: 500 V/us)

Freq (MHz) Vrms (mV) Required Slew Rate (V/us) Fraction of Slew Rate Used (%)
37
23.205
0.8586
0.1717
74
3.248
0.2403
0.0481
111
0.495
0.055
0.011
148
0.301
0.0445
0.0089
185
0.512
0.0947
0.0189

These calculations at least show that MAX 4107 should be much far away from reaching its slew rate limit in both RFPDs.

  2477   Mon Nov 11 20:42:49 2019 anchalDailyProgressFSSFSS Diagnostics - Two Tone Test

Method

  • RFPD of the FSS loops were tested through there Test input ports (Updated schematics: South SN010, North SN009)
  • I first took transfer function from the Test input port to the RF out port. Ideally, the transfer function of RFPD should be 100kOhm times this transfer function.
  • Moku is used to generate two RF signals with a different frequency equivalent to RFPD's resonant frequency.
  • These two outputs are combined with ZFSC-2-1W-S+ two-way splitter/combiner.
  • I switched around the voltage amplitude of these frequencies and took a spectrum from AG4395A near difference frequency.
  • The configuration file for spectrum measurement is present with data.
  • This measurement was made through twoToneTest script present in the data folder.

Inferences:

  • In first plot of each measurement set, I simply plotted the measured spectrum at the different signal voltage levels.
  • In second plots, I have done some back calculations, but I'm not sure if this is a good of estimating non-linearity.
  • I have used following formula to calculate ratio of total power in the difference frequency to the GM of powers of probe signals.
                                                                 \LARGE Ratio = \frac{ ASD_{@f2-f2} \sqrt{IFBW} } { \sqrt{ V_1\ TF_{@f1} V_2\ TF_{@f2}}}
  • This ratio I have plot in dB in the second plot.
  • It seems that the spectrum level generated is not dependent on how strong the probe signals are.
  • This can be seen as along the diagonal, instead of remaining the same ratio, the ratio decreases by 20 dB in each step which is an increase in the denominator above.
  • Lastly, I just connected the two ends of cable between which DUT goes and repeated the measurement.
  • The non-linearity of the measurement setup itself is safely 50 dB below the measured values.

But what now?

  • I need more time to make sense of these measurements. So that will come tomorrow.

Data

Attachment 1: NFSS_RFPD_SN009_TF_11-11-2019_183131.pdf
NFSS_RFPD_SN009_TF_11-11-2019_183131.pdf
Attachment 2: SFSS_RFPD_SN010_TF_11-11-2019_203310.pdf
SFSS_RFPD_SN010_TF_11-11-2019_203310.pdf
Attachment 3: FSStwoToneTest.pdf
FSStwoToneTest.pdf FSStwoToneTest.pdf FSStwoToneTest.pdf FSStwoToneTest.pdf FSStwoToneTest.pdf FSStwoToneTest.pdf FSStwoToneTest.pdf FSStwoToneTest.pdf
  2476   Sat Nov 9 19:28:36 2019 anchalDailyProgressFSSFSS Diagnostics - Loop nonlinearity Test

Method

I've wrote this script, nonlinTF.py which controls a Marconi 2023A and SR785 together. Marconi is used to providing a carrier frequency which is mixed with the Source Out signal from SR785 before feeding into the TEST2 input port on D040105 of TTFSS boxes. Then OUT1 port on D040105 of TTFSS box is used to read back at channel two of SR785 (channel one being fed with a copy of the Source Out signal). So SR785 is effectively measuring any downconversion in the loop (due to some nonlinearity) from micing of CF-IF, CF and CF+IF probe signals injected into the loop. The effectively closed-loop transfer function between TEST2 and OUT1 should be G/(1+G), so this injected signal should not suffer any suppression, nor should it affect the locks. The locks were maintained without any problems during the whole measurement. The CF frequency was stepped by 100 kHz from 100kHz to 10 MHz and then by 1 Mhz upto 100 MHz.

Mixer ZX05-1LHW (level +13 dBm) was used for the mixing and IF peak voltage was set to 30 mV. The configuration of the measurement for the transfer function is present in the configuration file in the folder.

Results?

  • I had to upload plots in png as they were aggregated result of a lot of data, pdf would have been very heavy.
  • North Side looks good with only minor bumps in 60 Hz harmonics. I'm not even sure if this just came because of SR785.
  • The South side however looks very dramatic. A lot of stuff happening when carrier frequency was near 50 MHz.
  • But I'm not sure if this measurement even makes any sense. South side is better in performance then North side, this wasn't expected.
  • Also, the measurements are long-time measurements, so a lot of things could have changed between North and South (they were taken on different days too).
  • I'll take more verifying measurements near the suspected frequencies tomorrow.
  • I'm also thinking of taking spectrums instead of swept sine transfer functions next time.

Data

Attachment 1: NFSS_NonLinearity_Test.png
NFSS_NonLinearity_Test.png
Attachment 2: SFSS_NonLinearity_Test.png
SFSS_NonLinearity_Test.png
Attachment 3: signal-attachment-2019-11-09-195635.jpeg
signal-attachment-2019-11-09-195635.jpeg
  2475   Sat Nov 9 13:46:08 2019 shrutiMiscElectronics EquipmentBorrowed

I've borrowed a Marconi VCO, a mini-circuits LPF and mixer from the CTN lab for use at the WB EEshop.


AG Mon Dec 30 10:24:28 2019 : Has been returned to CTN now.

  2474   Tue Nov 5 18:37:59 2019 anchalDailyProgressFSSFSS Diagnostics - TTFSS Testpoint Spectrum Data

Measurement method

  • All measurements were taken withAG4395A for high frequency and SR785 for low frequency and the corresponding measurement parameter files are attached.
  • For SR785, the SRmeasureWideSP.py script was used.
  • HO 41800A Active Probes were used with AG4395A.
  • For SR785, I simply used clips as it already has 1 MOhm input impedance. However, I have noticed oscillations in the spectrum.
  • Note, South FSS box had higher gains, Common Gain: 24 dB, Fast Gain: 14 dB, while the Northside had Common Gain: 10 dB and Fast Gain 10 dB.
  • TP1,5 and 4 are on common path on D040105 (North, South).
  • TP17,14,15,16,18 and 19 are on PZT Path on D040105 (North, South).
  • TP 11, 12 and 13 are on EOM Path on D040105 (North, South).

Data

Attachment 1: FSS_TP_Spectrum.pdf
FSS_TP_Spectrum.pdf FSS_TP_Spectrum.pdf FSS_TP_Spectrum.pdf FSS_TP_Spectrum.pdf FSS_TP_Spectrum.pdf FSS_TP_Spectrum.pdf FSS_TP_Spectrum.pdf FSS_TP_Spectrum.pdf FSS_TP_Spectrum.pdf FSS_TP_Spectrum.pdf FSS_TP_Spectrum.pdf FSS_TP_Spectrum.pdf
  2473   Mon Nov 4 20:13:44 2019 anchalNotesBEATQuick note: FSS Loop Gain Changes

South Common Gain: 24 dB! , Fast Gain: 14 dB

North Common Gain: 10 dB, Fast Gain: 10 dB

  2472   Mon Nov 4 15:03:26 2019 anchalDailyProgressFSSRaised North TTFSS; Fixed the boxes.

I raised the North TTFSS box by 6 inches to make way for working on South box and to reduce the congestion of connectors in front of the two boxes. I have also clamped the boxes to a fixed position now, so they can't move. This would ensure the cables are not hitting the end of the platform and face any severe strain.


The next step towards improving lab cable hygiene and layout is to replace all RF cables with RG-405 Belden-N 1671J cables. However, the effects of this change would be less significant then fixing the sick FSS. So I'll first focus on that.

Attachment 1: signal-attachment-2019-11-04-161746.jpeg
signal-attachment-2019-11-04-161746.jpeg
  2471   Mon Nov 4 12:15:01 2019 anchalNotesPMCChanging the autolocking method

Trying gain sliding like 40m

Rana told me that in 40m, the PMCs are autolocked by reducing the gain of the loop and changing the ramp until the lock is acquired. Then the gain is increased back to operation point. I tried this method with our South PMC as the usual method being used of 'changing Blanking state' wasn't working anymore. However, even with the gain set all the way to -10 dB, the loop was not locking exactly at the center of the TEM00 mode. And was unable to skip higher-order modes. There is a header H1 which changes the input stage gain. Removing this header pin, reducing the input stage gain by a factor of 100. Even after doing this, I was unable to robustly acquire the lock by this method. Besides, this reduced gain was the case earlier (CTN/2427) and it was too low as the VCA U5A AD602 had to be kept at maximum 30dB gain. So I did not want to reduce this first stage gain.


New method: Switch off input while changing Ramp

Somewhat similar to our FSS loops, I find it much cleaner to just not close the loop until we have reached near the lock point. This could be done fairly easily with the existing code. I just had to change the loopStateEnable variable from Engage (which changes the Blanking pin on U5A AD602) to input switch (FP1TEST for South and FP2TEST for North). So now, when finding a lock point, the input is changed to terminated inputs and the loop is closed when lock point is found. This works very nicely, just like the FSS autolocks.

This has finally fixed any problems with PMC autolocks.\LARGE {\color{Green} \checkmark}

  2470   Fri Nov 1 17:54:47 2019 anchalDailyProgressFSSFSS Diagnostics - RFPD RF Ouput under inspection

First step in FSS Diagnostics was to see RF output from the RFPDs in FSS when they are locked. I ran some extensive measurements to cover all the information about this signal. The RF out is sent to the FSS box through ZFDC-20-5-S+ 19.5 dB directional coupler. The coupled output's spectrum is measured at different frequency ranges using both AG4395A and SR785. The measurement configuration files are included with data for metadata of the measurement. The signal is also analyzed in time series with measurement upto 1 GSa/s with TDS 3034C and one measurement at 5 GSa/s with TDS 3052B. Both measurements were done manually setting minimum possible voltage resolution and using DC coupling with 50 Ohm impedance. All data is attached raw here for now. More interpretation and analysis to come soon.


Data

Attachment 1: NFSS_RFPD_Output_Oscilloscope.pdf
NFSS_RFPD_Output_Oscilloscope.pdf NFSS_RFPD_Output_Oscilloscope.pdf NFSS_RFPD_Output_Oscilloscope.pdf NFSS_RFPD_Output_Oscilloscope.pdf NFSS_RFPD_Output_Oscilloscope.pdf NFSS_RFPD_Output_Oscilloscope.pdf NFSS_RFPD_Output_Oscilloscope.pdf NFSS_RFPD_Output_Oscilloscope.pdf
Attachment 2: SFSS_RFPD_Output_Oscilloscope.pdf
SFSS_RFPD_Output_Oscilloscope.pdf SFSS_RFPD_Output_Oscilloscope.pdf SFSS_RFPD_Output_Oscilloscope.pdf SFSS_RFPD_Output_Oscilloscope.pdf SFSS_RFPD_Output_Oscilloscope.pdf SFSS_RFPD_Output_Oscilloscope.pdf SFSS_RFPD_Output_Oscilloscope.pdf SFSS_RFPD_Output_Oscilloscope.pdf
Attachment 3: NFSS_RFPD_RF_OUT_COUPLED.pdf
NFSS_RFPD_RF_OUT_COUPLED.pdf NFSS_RFPD_RF_OUT_COUPLED.pdf NFSS_RFPD_RF_OUT_COUPLED.pdf NFSS_RFPD_RF_OUT_COUPLED.pdf NFSS_RFPD_RF_OUT_COUPLED.pdf NFSS_RFPD_RF_OUT_COUPLED.pdf NFSS_RFPD_RF_OUT_COUPLED.pdf NFSS_RFPD_RF_OUT_COUPLED.pdf
Attachment 4: SFSS_RFPD_RF_OUT_COUPLED.pdf
SFSS_RFPD_RF_OUT_COUPLED.pdf SFSS_RFPD_RF_OUT_COUPLED.pdf SFSS_RFPD_RF_OUT_COUPLED.pdf SFSS_RFPD_RF_OUT_COUPLED.pdf SFSS_RFPD_RF_OUT_COUPLED.pdf SFSS_RFPD_RF_OUT_COUPLED.pdf SFSS_RFPD_RF_OUT_COUPLED.pdf SFSS_RFPD_RF_OUT_COUPLED.pdf
  2469   Wed Oct 30 18:35:31 2019 anchalNotesPMCIssue closed

It turns out I did not have a full understanding of the problem and it was not really a problem. The blanking (pin 4) on U5A AD602 doesn't shut down the channel, it just reduced the gain by 100. So if the gain in previous stage is large enough, the lock can still be acquired. And that's what was happening.

Ideally, we need to keep the AD602 on all the time and lock by scanning the offset with low gain. The loop will catch the lock (the exact same thing I thought was a problem) and once that has happened, we can just increase the loop gain the set value.


Presently, the gain behind the U5A AD602 is 101, which is kind of high. I just need to check if the above-mentioned locking method would work robustly without wrongly getting locked to any higher-order modes with the gain slider set to some threshold value within -10 dB to 30 dB. If that can't be done, I might have to reduce the gain in the first stage. For now, the cavities are locked and beatnote is traveling towards set point.

  2468   Wed Oct 30 15:15:28 2019 anchalNotesPMCReplacing U5A AD602 didn't solve the issue.

I brought a new AD602AR from 40m and replaced the U5A AD602 which from the previous post seemed like the culprit, but it wasn't. frown

I'll think of some new way of figuring out the point of the problem. It would be nice if someone can help me with this. All the history of the issue is on this thread starting at CTN:2451.


Edit Thu Oct 31 10:26:50 2019

Issue fixed. See CTN:2469.

  2467   Tue Oct 29 19:10:59 2019 anchalNotesPMCSouth PMC problem debugging efforts

I reduced the power falling on the PMC to ensure the high signal level isn't causing this problem. It was not. The problem still persisted.

Then, I did this reproducible step (quoted below)again, but this time I had a small 10 mV signal from SR785 going into FP2TEST and I was taking transfer function to TP2. If the U5A AD602 is switched off by the Blanking pin, the transfer function should remain null. This gave me a way of checking if the AD602 is wrongly getting switched on on its own.

  • To start, I kept Engage OFF. This gave a voltage of 4.41 V at pin 4 of U5A AD602. So it should be shut off.
  • The PZT voltage was about 97 V at this point.
  • The transfer function was flat to about -80 dB from 10Hz to 100Hz.
  • Then I started scanning the RAMP voltage of PZT. As the PZT voltage reached near 70V value, the PMC got locked on its own.
  • The transfer function value jumped suddenly to between -20 dB to 0 dB (an increase by about 70 dB).
  • The gain on AD602 was set to 0dB. So if it got on its own, the transfer function was expected to be 0dB.
  • The gate voltage at pin 4 of U5A AD602 was still 4.41 V. So ideally, it should still be off.
  • I have attached the data I captured. During swept sine, the ramp was being increased and we see a jump clearly when the PMC got locked on its own.

This is good evidence in my opinion that the AD602 at U5A is faulty. I need comments on this conclusion. If I don't hear otherwise by tomorrow noon, I'll start working on replacing it.


Data


Edit Thu Oct 31 10:26:50 2019

Issue fixed. See CTN:2469.

Quote:
 
  • But, when I connected the RFPD back, the problem was not there anymore. At this point, I found the following reproducible issue:
    • Engage button is off.
    • PZT displacement is following the ramp normally.
    • Slowly change the ramp and at the sweet spot, the PMC gets locked. Remember the engage is still off.
    • Now, changing the ramp doesn't change the PZT displacement anymore.
    • Switching on the engage ON doesn't change anything.
    • Switching it back OFF unlocks the PMC and the PZT displacement starts responding to ramp voltage again.
  • That's weird right. Since this was reproducible, I did this a few times and found that the problem doesn't necessarily happen at the locking point. It can happen anywhere. And in that case, the 5th step of switching on the engage does show a difference in locked mode. And again, switching it back OFF resolves the issue.
Attachment 1: SPMC_Diagnostics.pdf
SPMC_Diagnostics.pdf
  2466   Fri Oct 25 16:50:12 2019 anchalNotesPMCSouth PMC problem debugging efforts

Today, this problem happened again (check history for details). I have done the following investigative steps:

  • Situation: The South PMC PZT Ramp was changing by the EPICS voltage but the PZT displacement value was stuck. Also, the South Laser Slow voltage channel also goes back to some value even after changing it.
  • I took out the South PMC box from the rack. During this:
    • I had to disconnect the 9V supply used by acromag DACs for setting gain values and switch values.
    • I also had to disconnect the ethernet cable too.
    • I also disconnected RFPD input to the box, HV input, and PMC PZT.
  • The first thing I noticed is that the PMC PZT not following the ramp problem is fixed.
  • However, the slow output of the laser was going back to a set value (-6.1046V) without any reason.
  • I attached the SPMC RFPD RF (Downconverted) signal back to FPTest1 port the PMC just got locked! This happened even though the engage button was switched off.
    • And again, the PMC PZT ramp stopped causing any changes to the PZT displacement value. This is common when the PMC is locked in a normal manner.
  • I detached the SPMC RFPD, the changes were happening normally. So the presence of a strong signal might be the issue.
  • But, when I connected the RFPD back, the problem was not there anymore. At this point, I found the following reproducible issue:
    • Engage button is off.
    • PZT displacement is following the ramp normally.
    • Slowly change the ramp and at the sweet spot, the PMC gets locked. Remember the engage is still off.
    • Now, changing the ramp doesn't change the PZT displacement anymore.
    • Switching on the engage ON doesn't change anything.
    • Switching it back OFF unlocks the PMC and the PZT displacement starts responding to ramp voltage again.
  • That's weird right. Since this was reproducible, I did this a few times and found that the problem doesn't necessarily happen at the locking point. It can happen anywhere. And in that case, the 5th step of switching on the engage does show a difference in locked mode. And again, switching it back OFF resolves the issue.
  • At this point, I checked back the slow voltage and it had changed to -6.1086. Now apparently this was the new point it liked to stay on regardless of any changes made to the EPICS channel controlling it.
  • I checked if the signal for 'PMC Engage' to the U5A AD602. This signal works as expected. The gain setting signal also works as expected. The Ramp setting signal also works as expected. So there is no evidence that the acromag card or the EPICS channel is causing this issue (at least in the PMC problem).
  • It could be that the AD602 switches ON itself due to some reason and doesn't switch off until it explicitly sees a low-high change on its pin 4.
  • Also, it could be that the U9 PA85 is the one causing trouble. But I doubt that as the behavior matches more with engaging the lock upstream at U5A AD602.
  • I restarted the lab by restarting the channels and scripts (this is done in a preset protocol by restartAll command).
  • This resolved the issue with the slow voltage of the laser. But the PMC problem persists. In fact, it is not locking now.

I'll take another inspection with a fresh mind next time. This problem needs to be resolved as we can't leave some unexplained phenomenon to keep happening in the lab.

  2465   Thu Oct 24 11:42:06 2019 Ian MacMillanNotesFSSFSS Plant Model v2

With discussions with Anchal and some reading

It seems impossible to create a casual circuit with a zero phase shift. (See this for more) 

If we have a circuit with an impulse response h(t) and transfer function H(f)=F[h(t)] where H(-f)=H*(f). For the filter to cause no phase shift then ∠H(f)=0 for a complex exponential input for all f. It is also impossible to have a constant phase shift unless that phase shift is zero.

Therefore "filter does not change the phase at all, then H(f) is a real-valued function, and because of the conjugacy constraint, it is also an even function of f. But then its Fourier transform h(t) is an even function of time, and thus the filter cannot be causal (except in trivial cases): if its impulse response is nonzero for any particular t>0, then it is also nonzero for −t (where −t<0)" 

Because this can't be done casually, it should be done using a Field Programmable Gate Array. Unfortunately, I don't think we have access to one. I am reading up on the Moku FIR Filter builder to find out if we can program it to do what we want.

Quote:

I have designed a passive circuit that seems to match the ideal transfer functions in shape. Scaling should just be a game of playing with the values of the resistors and capacitors. The phase still seems to be an issue. There is an unwanted phase shift from 0 -> -90.

The next step is trying to finalize the values for the resistors and caps. Possibly model in zero if I have time. Then build and test. Also fix the phase.

Quote:

I have updated the plant model to contain the cavity pole also. Cavity pole is a pair of positive and negative real poles, so it is hard (or maybe impossible) to imitate it exactly with an electronic circuit. Or maybe, my analysis is wrong.

Nevertheless, I have for now made this circuit which has a second-order pole, so it correctly matches the magnitude of the model transfer function up to 1 MHz for both PZT and EOM paths. Note that the elliptical filter is not included in this as we can connect the circuit to Test port 1 which injects just before the filter in LIGO-D0901894. Also, for the gains in EOM path, I had to add some factors to make it the same as the model transfer function. All components are calculated for E12 series resistors and capacitors.

Attached is a pdf of the notebook which contains all the mathematics in latex and a zip file with all files to recreate and further work on this. Ian can use these as support to learn zero further.

 

 

  2464   Wed Oct 23 17:39:44 2019 anchalNotesPMCSouth PMC started working somehow :(

The same thing happened again. This time, not just with the SPMC actuation voltage, but the South Laser slow voltage control was also unresponsive. However, I am not very sure about the latter. This was resolved once the restarted the whole lab. This narrows down the problem to following possibilities:

  • One or more channels in an acromag card become unresponsive. It could be the 10.0.1.49 acromag used for South PMC Controls.
  • Some device in TCS or QIL lab is interfering by using an IP address reserved for docker services on which the EPICS channels are hosted and python scripts are run.

These still don't explain CTN:2456. Again, since this is an irreproducible error, I will just have to wait for it to happen to gather more clues. Right now, everything is fine and beatnote is traveling towards set frequency.

 

  2463   Wed Oct 23 13:21:12 2019 anchalDailyProgressNoiseBudgetFixed all discrepancies.

I ran the code for aLIGO coating structure and tried to reproduce fig5 and fig7 of Hong et al. paper. It turned out that the derivatives of the complex reflectivity were not matching with the paper. I rewrote the code, with a fresh mind without looking at previous code and voila, after increasing computation time slightly due to more brute force calculations, I was able to reproduce the figures. These figures are attached. Since I do not have access to the data of the figures in the paper (I tried to email the authors but got no replies), I could only try to plot it on the same scale and limits and check the values by eyes. The values seem to match. So I am more confident now to declare that this code completely follows the paper's calculations.


However, this does not change the coating Brownian noise. I have updated the noise budgets at the Daily and Latest plots.

Attachment 1: aLIGOComplexReflectivityVariationFig5.pdf
aLIGOComplexReflectivityVariationFig5.pdf aLIGOComplexReflectivityVariationFig5.pdf aLIGOComplexReflectivityVariationFig5.pdf aLIGOComplexReflectivityVariationFig5.pdf
Attachment 2: aLIGOBreakDownThermalNoiseFig7.pdf
aLIGOBreakDownThermalNoiseFig7.pdf aLIGOBreakDownThermalNoiseFig7.pdf aLIGOBreakDownThermalNoiseFig7.pdf aLIGOBreakDownThermalNoiseFig7.pdf
Attachment 3: aLIGOCoatingBrNoiseCalculations.zip
  2462   Tue Oct 22 13:07:45 2019 Ian MacMillanNotesFSSFSS Plant Model v2

I have designed a passive circuit that seems to match the ideal transfer functions in shape. Scaling should just be a game of playing with the values of the resistors and capacitors. The phase still seems to be an issue. There is an unwanted phase shift from 0 -> -90.

The next step is trying to finalize the values for the resistors and caps. Possibly model in zero if I have time. Then build and test. Also fix the phase.

Quote:

I have updated the plant model to contain the cavity pole also. Cavity pole is a pair of positive and negative real poles, so it is hard (or maybe impossible) to imitate it exactly with an electronic circuit. Or maybe, my analysis is wrong.

Nevertheless, I have for now made this circuit which has a second-order pole, so it correctly matches the magnitude of the model transfer function up to 1 MHz for both PZT and EOM paths. Note that the elliptical filter is not included in this as we can connect the circuit to Test port 1 which injects just before the filter in LIGO-D0901894. Also, for the gains in EOM path, I had to add some factors to make it the same as the model transfer function. All components are calculated for E12 series resistors and capacitors.

Attached is a pdf of the notebook which contains all the mathematics in latex and a zip file with all files to recreate and further work on this. Ian can use these as support to learn zero further.

 

Attachment 1: PlantDig.pdf
PlantDig.pdf
Attachment 2: EOM_Transfer_Function.pdf
EOM_Transfer_Function.pdf
Attachment 3: PZT_Transfer_Function.pdf
PZT_Transfer_Function.pdf
  2461   Fri Oct 18 17:47:49 2019 anchalNotesBEATQuick note: BN Noise bakc to where it was.

Latest BN Spectrum: CTN_Latest_BN_Spec.pdf

Daily BN Spectrum: CTN_Daily_BN_Spec.pdf


  • Changed bandwidth of Moku PLL 10 10 kHz until beatnote frequency control settles back normally.
  • We are back to same position in noise but there are some extra sharp peaks which need to be looked at. But they are very small.
  2460   Wed Oct 16 19:11:11 2019 anchalNotesBEATQuick note: Cavities locked; Changed BN Freq Counter Poll Rate
  • Follow discussion in thread  CTN:2451.
  • Changed polling rate of beatnote frequency counter to 1 Hz and the script is updating the EPICS channel at the same rate now.
  • Changed BNFreqSignUpdate and BNFreqSlopeCalc scripts accordingly.
ELOG V3.1.3-